OpenStudy (daniellelovee):

I have two machines a pulley and a block and tackle I have 5 weights weighing 1 gram 2 grams 3 grams 4 grams and 5 grams. On each machine I raise the weight 5 cm. pulley 1 gram = 1 newtons of downward force 2 gram = 2 newtons of downward force 3g = 3n 4g = 4n 5g = 5n Block and tackle 1g = .5n 2g = 1n 3g = 1.5n 4g = 2n 5g = 2.5n

OpenStudy (daniellelovee):

How did the work output relate to the work input for the pulleys? Explain using your data.

OpenStudy (daniellelovee):

@Michele_Laino @Somy

OpenStudy (michele_laino):

from the data, I get this: work done by the pulley, is \(L_P= F \cdot d\) whereas, work done by the block and tackle, is \(L_{B-T}=(F/2) \cdot d\) where \(d=5\) cm so we have this ratio: \[\frac{{{L_P}}}{{{L_{B - T}}}} = \frac{{F \cdot d}}{{\left( {F/2} \right) \cdot d}} = ...?\]

OpenStudy (daniellelovee):